Give asymptotic upper and lower bounds for T(n) in each of the following recurrences. Assume that T(n) is constant for n ≤ 3. Make your bounds as tight as possible, and justify your answers.

5.a

T(n) = 2T(n/3) + n lg n

5.b

T(n) = 7T(n/2) + n3

5.c

T(n) = 3T(n/5) + lg2 n
Answer #1

Answer

a)
T(n) = 2T(n/3)+n\lg{n}\\ This\; is \;in \;the \;form\;of \;T(n) = aT(\frac{n}{b}) + f(n)\\ so,\;we\;can\;use\;master's\;theorem.\\ compare\;n^{\log_b{a}}\;with\;f(n)\\ a=2,\;b=3\;\\ n^{\log_b{a}}=n^{\log_3{2}}=n^{0.64}\\ n^{\log_b{a}}=n^{\log_3{2}}=n^{0.64}\;<\;f(n)=n\lg{n}\\ so,\;this\;comes\;under\;case\;3\\ so,\;time\;complexity\;is\;\Theta(f(n))=\Theta(n\lg{n})\\

b)
T(n) = 7T(n/2) + n^{3}\\ This\; is \;in \;the \;form\;of \;T(n) = aT(\frac{n}{b}) + f(n)\\ so,\;we\;can\;use\;master's\;theorem.\\ compare\;n^{\log_b{a}}\;with\;f(n)\\ a=7,\;b=2\;\\ n^{\log_b{a}}=n^{\log_2{7}}=n^{2.8}\\ n^{\log_b{a}}=n^{\log_2{7}}=n^{2.8}\;<\;f(n)=O(n^{3})\\ so,\;this\;comes\;under\;case\;3\\ so,\;time\;complexity\;is\;\Theta(n^{3})\\

c)
T(n) = 3T(n/5) + \lg^2{n}\\ This\; is \;in \;the \;form\;of \;T(n) = aT(\frac{n}{b}) + f(n)\\ so,\;we\;can\;use\;master's\;theorem.\\ compare\;n^{\log_b{a}}\;with\;f(n)\\ a=3,\;b=5\;\\ n^{\log_b{a}}=n^{\log_5{3}}=n^{0.69}\\ n^{\log_b{a}}=n^{\log_5{3}}=n^{0.69}\;>\;f(n)=\lg^2{n}\\ so,\;this\;comes\;under\;case\;1\\ so,\;time\;complexity\;is\;\Theta(n^{\log_5{3}})=\Theta(n^{0.69})\\